# Predicting Networks Before They Happen: Experimentation on a Real-Time V2X Digital Twin

### v1

@@ III-B Computational Complexity and Execution Time
Let $\tau_{e2e}$ be the total time elapsed between the acquisition of a measurement by a Physical Entity and the completion of the corresponding DT update. Formally:
@@
where $\tau_{m}$ is the measurement acquisition time, $\tau_{w}$ the communication delay between vehicles and the Mobility DT, $\tau_{tp}$ the Mobility DT update and trajectory prediction time, $\tau_{req}$ the inter-process communication delay, and $\tau_{rt}$ the execution time of the physical-modeling stage. Given the computational complexity of the ray tracing operation, $\tau_{rt}$ is the dominant and most variable value. Specifically, it depends on the level of physical detail included in the propagation model. To describe these configurations in a compact and general manner, we introduce the Ray Tracing Detail Index, denoted DI, a scalar metric used to classify the fidelity level of the physical-modeling stage. Each DI level corresponds to a specific configuration of the physical model, ranging from lightweight approximations to high-fidelity reconstructions. Table I reports the five DI levels considered in this work. Lower levels include only the essential components needed for coarse propagation estimates, whereas higher levels progressively incorporate more detailed interaction mechanisms, finer spatial resolution, and deeper multipath exploration. Although in our implementation these fidelity levels map to different configurations of a ray-based propagation model, the abstraction provided by DI remains general and applicable to any physical-modeling backend supporting multiple complexity tiers. As the DI increases, the execution time of the physical model, denoted $\tau_{rt}$, grows accordingly. To ensure causal consistency, the prediction horizon $h$ must exceed the end-to-end latency $\tau_{e2e}$. Increasing model fidelity therefore requires larger horizons, which in turn amplify trajectory prediction errors and their impact on channel estimation, revealing a fundamental trade-off between physical-model complexity, latency, and predictive accuracy. Section IV provides an empirical characterization of these trade-offs based on real-world experiments conducted in Tokyo, Japan, evaluating how different DI levels impact both computational latency and channel prediction accuracy.
@@ IV Experimental and Simulated Results
To demonstrate the practical operation of the proposed system, we consider a proof-of-concept deployment carried out in the Ookayama campus of the Institute of Science Tokyo, Japan complete with measured end-to-end latencies. To complement the experimental analysis, we then consider a simulation study performed in the identical geometric scenario in which we systematically vary the DI and introduce synthetic perturbations into the trajectory prediction module.
@@ IV-A Scenario introduction
Buildings and Roads Material Concrete [iturp2040]
Trees Material Wood [iturp2040]
Cars Material Metal [iturp2040]
@@
The evaluation scenario involves three connected vehicles (Figure 3(a)) that interact with the Mobility DT through a $2.4$ GHz Wi-Fi CP, and are each equipped with a roof-mounted Panasonic WiGig RSU forming the $60$ GHz User Plane (UP). The Mobility DT and the VaN3Twin engine are deployed at the edge—corresponding to the location of Vehicle 3—and exchange data through a wired Ethernet LAN. Figure 3(b) illustrates the corresponding VaN3Twin ray tracing scenario of the considered street in Ookayama, Tokyo, Japan. Buildings, vehicles, and antennas are modeled in VaN3Twin using material properties and radiation patterns consistent with the experimental setup (see Table II). VaN3Twin ray tracing details are reported in Table II. During the experimental campaign, the vehicles perform iperf-based data transmissions over the WiGig UP. Vehicle 1 (a Toyota COMS) and Vehicle 3 (a motorized wheel cart) remain stationary and maintain a direct Line of Sight (LoS) link, while Vehicle 2 (a Toyota Estima) follows a predefined trajectory (Figure 3(b)) that intermittently obstructs the link. The Mobility DT is used to predict, in real time, the $60$ GHz channel between Vehicles 1 and 3 via VaN3Twin, based on the predicted trajectory of Vehicle 2 with prediction horizon $h=500$ ms. The Mobility DT simultaneously collects from each vehicle its current position and the field-measured Received Signal Strength Indicators (RSSIs), which are employed for in-the-loop calibration of the prediction process. After computation, the predicted RSSI values are transmitted to the vehicles through the Wi-Fi CP. To quantify the effectiveness of this real-time predictive loop, we evaluate the accuracy of the RSSI forecasts produced by VaN3Twin along the trajectory of Vehicle 2. The Mobility DT employs the simple predictor for short-term trajectory estimation presented in Section III-A, while VaN3Twin operates using the DI 5 configuration reported in Table I. Results show a maximum average RSSI prediction error of $1.01$ dB and a 95th-percentile error of $3.05$ dB, demonstrating accurate prediction of the $60$ GHz link dynamics under mobility-induced obstruction. Since VaN3Twin operates on predicted positions, channel accuracy is jointly limited by trajectory error and by the ray tracing DI, which directly determines the end-to-end latency and thus the minimum prediction horizon required for real-time operation.
@@ IV-B End-to-End System Latency
Building on the formal definition introduced in Section III-B, this section reports on-field measurements of the end-to-end latency $\tau_{\mathrm{e2e}}$, decomposed according to (4). All the measured fixed latency components are summarized in Table III, while the resulting end-to-end latency distributions as a function of the DI level are shown in Figure 4.
Fixed system latency components: measurement acquisition ($\tau_{m}$) was evaluated by collecting location and channel samples over $60$ s with a $100$ ms sampling interval, yielding approximately $600$ samples per experiment. No statistically significant differences were observed between static and dynamic conditions. Communication-related delays ($\tau_{w}$ and $\tau_{req}$) were measured using ICMP-based Round Trip Time (RTT) probing over $100$ messages and conservatively converted to one-way latency by halving the resulting value. Mobility DT update time and trajectory prediction latency ($\tau_{tp}$) was obtained by profiling the time from the scenario update up to the execution of the linear prediction algorithm described in Section III-A. These components—resumed in Table III—constitute a fixed system latency budget, independent of the DI configuration.
@@
End-to-end latency $\tau_{\mathrm{e2e}}$: is obtained by combining the fixed system latency components with the DI-dependent ray tracing execution time. The resulting distributions of $\tau_{\mathrm{e2e}}$ for each DI configuration are reported in Figure 4, highlighting the progressive increase in both latency and variability as the ray tracing complexity grows. These results demonstrate that improving channel modeling fidelity by increasing the DI level inevitably enlarges the end-to-end latency budget. Since channel prediction must be available before transmission, this latency directly translates into a minimum required trajectory prediction horizon. As a result, higher-fidelity channel modeling forces the DT to rely on longer-term mobility forecasts.
@@ IV-C Impact of trajectory prediction error
Because the end-to-end latency of the DT determines how far into the future vehicle positions must be predicted, longer latencies inevitably amplify trajectory prediction errors. This section quantifies how such latency-induced positional inaccuracies propagate into ray-traced channel prediction.
Let $\mathbf{p}(t)=[x(t),\,y(t)]^{\top}$ denote the ground-truth position at time $t$, and let $\tilde{\mathbf{p}}(t)=\mathbf{p}(t)+\mathbf{e}(t)$, be its perturbed counterpart, where $\mathbf{e}(t)$ models the prediction error. A normalized parameter $k\in[0,1]$ controls the maximum error magnitude through $\varepsilon_{k}=k\,\varepsilon_{\max}$, with $\varepsilon_{\max}=1$ m in the considered setup. The displacement vector is sampled uniformly from a circular annulus $\mathbf{e}(t)\sim\mathcal{A}\!\left(\tfrac{1}{3}\varepsilon_{k},\;\tfrac{1}{2}\varepsilon_{k}\right),$ and $\mathbf{e}(t)$ is isotropically sampled within the annulus itself. To quantify the impact of trajectory prediction inaccuracies, we evaluate two complementary metrics: (i) the deviation in received signal strength and (ii) the stability of the geometric visibility state. Both metrics are computed by comparing the channels obtained from the perturbed positions $\tilde{\mathbf{p}}(t)$ with those obtained from the true positions $\mathbf{p}(t)$.
@@
where $N$ is the number of channel realizations. Figure 5 shows the dependence of the computed RMSE (blue) on the perturbation level $k$, focusing on measurements collected during the LoS to non-LoS transition caused by the obstructing vehicle (Vehicle 2). The relationship is clearly non-linear, with a rapid error increase for small positional deviations. For the considered scenario, $k\approx 0.1$ already yields an RMSE of about $20$ dB on RSSI prediction. Since small geometric inaccuracies already induce large prediction errors, these results also anticipate the sensitivity of the LoS classification, analyzed next.
@@ V Conclusions
In this work, we presented and experimentally validated a real-time, predictive NDT for V2X communications, explicitly designed around end-to-end latency constraints. By integrating a live Mobility Digital Twin with the VaN3Twin engine, we implemented a closed-loop workflow capable of operating in synchrony with the physical environment. A key outcome of this study is the end-to-end latency analysis of the Digital Twin pipeline, covering sensing, mobility prediction, ray-tracing-based channel computation, and data exchange. This analysis allowed us to identify how the computational complexity of the physical model directly constrains the feasible prediction horizon and, consequently, the usefulness of the Digital Twin for real-time operation. Experimental results obtained in a live urban environment demonstrate that high-fidelity channel modeling can be executed within a bounded time budget, provided that model complexity and scheduling are jointly designed. Overall, this work shows that real-time integration of mobility Digital Twins and ray-tracing-based network modeling is feasible and practically deployable. Future work will extend this latency-aware framework toward real-time prediction of higher-layer network performance metrics leveraging on VaN3Twin full-stack V2X simulation capability.
@@ Acknowledgment
This work was partially supported by the European Union - Next Generation EU under the Italian National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.3,CUP D43C22003080001, partnership on “Telecommunications of the Future” (PE00000001 - program “RESTART”).

### v2

@@ III-B Computational Complexity and Execution Time
Let $\tau_{e2e}$ be the total time elapsed between the acquisition of a measurement by a vehicle and the completion of the corresponding DT update. Formally:
@@
where $\tau_{m}$ is the measurement acquisition time, $\tau_{w}$ the communication delay between vehicles and the Mobility DT, $\tau_{tp}$ the Mobility DT update and trajectory prediction time, $\tau_{req}$ the inter-process communication delay, and $\tau_{rt}$ the execution time of the physical-modeling stage. Given the computational complexity of the ray tracing operation, $\tau_{rt}$ is the dominant and most variable component. Specifically, it depends on the level of physical detail included in the propagation model. To describe these configurations in a compact and general manner, we introduce the Ray Tracing Detail Index, denoted DI, a scalar metric used to classify the fidelity level of the physical-modeling stage. Each DI level corresponds to a specific configuration of the physical model, ranging from lightweight approximations to high-fidelity representations. Table I reports the five DI levels considered in this work. Lower levels include only the essential components needed for coarse propagation estimates, whereas higher levels progressively incorporate more detailed interaction mechanisms, finer spatial resolution, and higher-order multipath modeling. Although in our implementation these fidelity levels map to different configurations of a ray-based propagation model, the abstraction provided by DI remains general and applicable to any physical-modeling backend that supports multiple complexity tiers. As the DI increases, the execution time of the physical model, denoted $\tau_{rt}$, grows accordingly.
To ensure causality, the prediction horizon $h$ must exceed the end-to-end latency $\tau_{e2e}$. Increasing model fidelity therefore requires larger horizons, which in turn amplify trajectory prediction errors and their impact on channel estimation, revealing a fundamental trade-off between physical-model complexity, latency, and predictive accuracy. Section IV provides an empirical characterization of these trade-offs based on real-world experiments conducted in Tokyo, Japan, evaluating how different DI levels impact both computational latency and channel prediction accuracy.
@@ IV Experimental and Simulated Results
To demonstrate the practical operation of the proposed system, we consider a proof-of-concept deployment carried out in the Ookayama campus of the Institute of Science Tokyo, Japan including measured end-to-end latencies. To complement the experimental analysis, we then consider a simulation study performed in the identical geometric scenario in which we systematically vary the DI and introduce synthetic perturbations into the trajectory prediction module.
@@ IV-A Scenario introduction
Buildings and Roads Material Concrete [8]
Trees Material Wood [8]
Cars Material Metal [8]
@@
The evaluation scenario involves three connected vehicles (Figure 3(a)) that interact with the Mobility DT through a $2.4$ GHz Wi-Fi CP, and are each equipped with a roof-mounted Panasonic WiGig RSU forming the $60$ GHz high-capacity User Plane (UP). The Mobility DT and VaN3Twin are deployed at the edge—corresponding to the location of Vehicle 3—and exchange data through a wired Ethernet LAN. Figure 3(b) illustrates the corresponding VaN3Twin ray tracing scenario of the considered street in Ookayama, Tokyo, Japan. Buildings, vehicles, and antennas are modeled in VaN3Twin using material properties and radiation patterns consistent with the experimental setup (see Table II). VaN3Twin ray tracing details are reported in Table II. During the experimental campaign, the vehicles perform iperf-based data transmissions over the WiGig UP. Vehicle 1 (a Toyota COMS) and Vehicle 3 (a motorized wheel cart) remain stationary and maintain a direct Line of Sight (LoS) link, while Vehicle 2 (a Toyota Estima) follows a predefined trajectory (Figure 3(b)) that intermittently blocks the LoS path. The Mobility DT is used to predict, in real time, the $60$ GHz channel between Vehicles 1 and 3 via VaN3Twin, based on the predicted trajectory of Vehicle 2 with prediction horizon $h=500$ ms. Such predictions may be employed, for example, for proactive network topology management. The Mobility DT simultaneously collects from each vehicle its current position and the field-measured Received Signal Strength Indicators (RSSIs), which are employed for in-the-loop calibration of the prediction process. After computation, a control message is sent to the vehicles through the Wi-Fi CP. To quantify the effectiveness of this real-time predictive loop, we evaluate the accuracy of the RSSI forecasts produced by VaN3Twin along the trajectory of Vehicle 2. The Mobility DT employs the simple predictor for short-term trajectory estimation presented in Section III-A, while VaN3Twin operates using the DI 5 configuration reported in Table I. Results show a maximum average RSSI prediction error of $1.01$ dB and a 95th-percentile error of $3.05$ dB, demonstrating accurate prediction of the $60$ GHz link dynamics under mobility-induced obstruction. Since VaN3Twin operates on predicted positions, channel accuracy is jointly limited by trajectory prediction error and by the selected ray tracing DI, which directly determines the end-to-end latency and thus the minimum prediction horizon required for real-time operation.
@@ IV-B End-to-End System Latency
Building on the formal definition introduced in Section III-B, this section reports on-field measurements of the end-to-end latency $\tau_{\mathrm{e2e}}$, decomposed according to (4).
Fixed system latency components: are summarized in Table III and constitute a fixed system latency budget, independent of the DI configuration. Measurement acquisition ($\tau_{m}$) was evaluated by collecting location and channel samples over $60$ s with a $100$ ms sampling interval, yielding approximately $600$ samples per experiment. No statistically significant differences were observed between static and dynamic conditions. Communication-related delays ($\tau_{w}$ and $\tau_{req}$) were measured using ICMP-based Round Trip Time (RTT) probing over $100$ messages and conservatively converted to one-way latency by halving the resulting value. Mobility DT update time and trajectory prediction latency ($\tau_{tp}$) was obtained by profiling the time from the scenario update up to the execution of the linear prediction algorithm described in Section III-A.
@@
End-to-end latency $\tau_{\mathrm{e2e}}$: is obtained by combining the fixed system latency components with the DI-dependent ray tracing execution time. The distributions of $\tau_{\mathrm{e2e}}$ in Figure 4 show increasing latency and variability with ray tracing complexity, peaking at an average $250$ ms for DI 5.
These results demonstrate that improving channel modeling fidelity by increasing the DI level inevitably enlarges the end-to-end latency budget. Since channel prediction must be available before transmission, this latency directly translates into a minimum required trajectory prediction horizon. As a result, higher-fidelity channel modeling forces the DT to rely on longer-term mobility forecasts.
@@ IV-C Impact of trajectory prediction error
Because the end-to-end latency of the DT determines how far into the future vehicle positions must be predicted, longer latencies inevitably exacerbate trajectory prediction errors. This section quantifies how such latency-induced positional inaccuracies propagate into ray-traced channel prediction.
Let $\mathbf{p}(t)=[x(t),\,y(t)]^{\top}$ denote the ground-truth position at time $t$, and $\tilde{\mathbf{p}}(t)=\mathbf{p}(t)+\mathbf{e}(t)$ its perturbed counterpart, where $\mathbf{e}(t)$ models the prediction error. A normalized parameter $k\in[0,1]$ controls the maximum error magnitude through $\varepsilon_{k}=k\,\varepsilon_{\max}$, with $\varepsilon_{\max}=1$ m in the considered setup. The displacement vector is sampled uniformly from a circular annulus $\mathbf{e}(t)\sim\mathcal{A}\!\left(\tfrac{1}{3}\varepsilon_{k},\;\tfrac{1}{2}\varepsilon_{k}\right),$ and $\mathbf{e}(t)$ is isotropically sampled within the annulus itself. To quantify the impact of trajectory prediction inaccuracies, we evaluate two complementary metrics: (i) the deviation in received signal strength and (ii) the stability of the geometric visibility state. Both metrics are computed by comparing the channels obtained from the perturbed positions $\tilde{\mathbf{p}}(t)$ with those obtained from the true positions $\mathbf{p}(t)$.
@@
where $N$ is the number of channel realizations. Figure 5 shows the dependence of the computed RMSE (blue) on the perturbation level $k$, focusing on measurements collected during the LoS to non-LoS transition caused by the obstructing vehicle (Vehicle 2). The relationship is clearly non-linear, with a rapid increase in error for small positional deviations. For the considered scenario, $k\approx 0.1$ already yields an RMSE of approximately $20$ dB in the RSSI prediction. Since small geometric inaccuracies already induce large prediction errors, these results also foreshadow the sensitivity of the LoS classification, analyzed next.
@@ V Conclusions
In this work we presented and experimentally validated a real-time, predictive NDT for V2X communications, explicitly designed around end-to-end latency constraints. By integrating a live Mobility Digital Twin with the VaN3Twin engine, we implemented a closed-loop workflow capable of operating in synchrony with the physical environment. A key outcome of this study is the end-to-end latency analysis of the Digital Twin pipeline, covering sensing, mobility prediction, ray-tracing-based channel computation, and data exchange. This analysis allowed us to identify how the computational complexity of the physical model directly constrains the feasible prediction horizon and, consequently, the usefulness of the Digital Twin for real-time operation. Experimental results obtained in a live urban environment demonstrate that high-fidelity channel modeling can be executed within a bounded time budget, provided that model complexity and scheduling are jointly designed. Overall, this work shows that real-time integration of mobility Digital Twins and ray-tracing-based network modeling is feasible and practically deployable. Future work will extend this latency-aware framework toward real-time prediction of higher-layer network performance metrics leveraging VaN3Twin full-stack V2X simulation capability.
@@ Acknowledgment
This work was partially supported by the European Union - Next Generation EU under the Italian National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.3, CUP D43C22003080001, partnership on “Telecommunications of the Future” (PE00000001 - program “RESTART”). This work was also supported by Japan Science and Technology Agency (JST) as part of Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE), Grant Number JPMJAP2517.